OpenStudy (anonymous):

how do you convert a vector that has something like v=3x-1y into polar coordinates of v=a(r)+b(omega) ?

OpenStudy (anonymous):

x=rcos(theta) y=r sin(theta)

OpenStudy (anonymous):

also, r = root(x^2 + y^2)

OpenStudy (anonymous):

so, you equation would look something like \[3 r \cos (\theta )-r \sin (\theta )\]

OpenStudy (anonymous):

so when i do that equation, x=rcos(theta would i use 3 as my x?

OpenStudy (anonymous):

When you see x, replace it with r cos (theta) When you see y, replace it with r sin (theta)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

imran, one more question, you couldn't simplify v=2rcos(θ)-3rsin(θ) right? am i not thinking?

OpenStudy (anonymous):

you can factor out a 'r' but that does not make much difference r(2 Cos[x]-3 Sin[x])

OpenStudy (anonymous):

alrighty cool. thank you!